A petrol pump station has two pumps. The







service times follows the exponential







distribution with a mean of 4 minutes and







cars arrive for service in a Poisson process at







the rate 10 cars per hour. Find the







probability of that a customer has to wait for







service.

1
Expert's answer
2022-02-15T11:13:18-0500

 "\begin{gathered}

\mu=\frac{60}{4}=15 \text { per hour } \\

\lambda=10 \text { per hour }

\end{gathered}"

P(Customer has to wait for service is) "= \frac{1}{\mu-\lambda}"

"\begin{aligned}

&=\frac{1}{15-10} \\

&=0.2

\end{aligned}"

Now proportion of time pumps remain idle.

So, it can be explained by formula "\frac{\lambda}{\mu}."

"\text { i.e. } \frac{\lambda}{\mu}=\frac{10}{15}=0.66"